OpenStudy (anonymous):

(cosec x - sin x) (sec x - cos x) = 1 / tan x +cot x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

LHS: (1-sin^2x)/sinx * (1-cos^2x) /cosx ={cos^2xsin^2x}/sinxcosx =sinxcosx RHS: tanx/tan^2x+1 =tanx/sec^2x =sinxcosx Hence proved.

mathslover (mathslover):

\(\large{( \csc x - \sin x)(\sec x - \cos x) = \cfrac{1}{\tan x + \cot x}}\) [ RTP - Required to Prove] \((\cfrac{1}{\sin x} - \sin x)( \cfrac{1}{\cos x} - \cos x )\) \((\cfrac{1-\sin ^2 x}{\sin x} )(\cfrac{1- \cos^2 x}{\cos x})\)

mathslover (mathslover):

Use : \(1 - \sin^2 x = \cos ^2 x\) and \(1- \cos ^2 x = \sin^2 x\) .

OpenStudy (anonymous):

Just what I did, captain obvious. ^ -_-"

mathslover (mathslover):

\(\large{\cfrac{\sin x \cos x}{1} = \cfrac{\sin x \cos x}{\sin^2 x + \cos^2 x}}\) As : \(1 = \sin^2 x + \cos^2 x\) \(\large{\cfrac{\cfrac{\sin x \cos x}{\sin x \cos x}}{\cfrac{\sin^2 x}{\sin x \cos x} + \cfrac{\cos ^2 x}{\sin x \cos x}}}\) \(\large{\cfrac{1}{\cfrac{\cancel{\sin^2 x}^ {\sin x}}{\cancel{\sin x} \cos x} + \cfrac{\cancel{\cos^2 x }^{\cos x}}{\cancel{\cos x} \sin x}}}\) \(\large{\cfrac{1}{\tan x + \cot x}}\)

mathslover (mathslover):

@NeetziD now feel the difference. You solved the LHS first to get \(\sin x \cos x\) and then solved RHS to get the same : \(\sin x \cos x\) . Sorry, but I consider that as verifying not as proof , it is a proof but not a perfect one. There may be some other opinions regarding my thinking. Though, your method is 99% correct according to me. I just presented a method where you can use LHS to find RHS .

OpenStudy (raden):

agree ^^

mathslover (mathslover):

@kewal I hope you got it. \( \begin{array}{} \cal{\large{\text{Hey! Good to see you on this site!}}}\\ \\ {\huge \mathtt{\color{blue}{\text{Welcome To OpenStudy!}}}} \\ \text{enjoy answering and asking questions here}\\ \cal{\large{\text{Tips:)}}} \\ \textbf{Ambassadors (people with A next to their name) are here to help you.}\\ \textbf{If you have any questions about the site, feel free to ask us!}\\ \textbf{Users who appear in Purple are Moderators and help run the site.}\\ \textbf{If you have a serious issue with the site or with the behavior of other}\\ \textbf{users}\\ \textbf{please feel free to raise it with them.}\\ \textbf{Want to study a particular subject? Click 'Find More Subjects'}\\ \textbf{search for it there.}\\ \textbf{If you get a good response from a user , click on best response}\\ \end{array}\)

OpenStudy (anonymous):

get it... =)

OpenStudy (anonymous):

@mathsolver very much thnx for yur help, r u math teacher..?

mathslover (mathslover):

You're welcome Kewal and no, I am not a math teacher. Just a student like you :)

OpenStudy (anonymous):

i jst came 2 tenth grade this year and was having problems with trigo.. thanx again fr your help.

mathslover (mathslover):

Good, keep it up!